OpenStudy (cookie_2046):

Which answer is the best estimate of x based on rounding the constants in the equation to the nearest interger? 5.2x+3.2=9.5 A. 1.1 B. 1.5 C. 1.2 D. 1.4

OpenStudy (anonymous):

hmm the answer will be 1.2

OpenStudy (anonymous):

C?

OpenStudy (anonymous):

Work: 5.2x = 9.5 - 3.2 x = 6.3/5.2 x = 1.21133 rounding it off :) x = 1.2 the answer is c